OpenStudy (tyrinzoebrown):

Can someone help Solve the equation -5= s-3 Show all work

OpenStudy (debbieg):

Add 3 to both sides. That will give you s=?

OpenStudy (tyrinzoebrown):

like this? 3-5=s-3+3

OpenStudy (debbieg):

Sure, that will work.

OpenStudy (debbieg):

So s= ?

OpenStudy (tyrinzoebrown):

s=-2 ?

OpenStudy (debbieg):

Exactly :)

OpenStudy (anonymous):

-5= s-3 -5+3=s -2=s

OpenStudy (tyrinzoebrown):

Thank you!

OpenStudy (debbieg):

You're welcome :)